OpenStudy (anonymous):

A bag contains 3 white marbles, 2 black marbles, and 5 red marbles. Find the probability of picking a red marble? Answers: 1/2 3/10 1/5 1/4

OpenStudy (anonymous):

1/2

OpenStudy (aravindg):

3/3+2+5 =3/10

OpenStudy (aravindg):

fav/total=3/10

OpenStudy (aravindg):

oopts its 5/10 =1/2

OpenStudy (aravindg):

because red balls =5

OpenStudy (anonymous):

1/2

OpenStudy (aravindg):

ok next qn :)